# Exploratory Causal Inference in SAEnce

### v1

@@ Appendix B Neural Effect Test
2:  // A) Arm-wise residualize only the tested neuron $j$
@@ Appendix C Minimal Python Implementation Snippet
We provide a minimal Python snippet for our algorithm relying on pandas library [McKinney et al., 2011] for tabular operations and SciPy library for statistical testing [Virtanen et al., 2020] .
@@
By stratification with median binning. For simplicity we ignore here arm-wise residualization (see full code for detailed implementation).
@@ D.1 CelebA semi-synthetic RCTs
Dataset. We use CelebA [Liu et al., 2018], a face attributes dataset with $200\mathrm{k}$ images and $40$ binary attributes per image  [^2] It can be downloaded from flwrlabs/celeba. Furthermore, for implementation details, labels have been doubled (we pass from Beard to Has_Beard and Has_notBeard). We follow the authors’ official train/val/test split, and we employ the validation data for training SAEs and the test data to interpret them. Attributes are treated as ground-truth binary labels. From this source, we simulate several RCTs following the data generating process (DGP) described below, varying the sample size ($n\ll 200k$) and treatment effect ($\tau$), reflecting realistic randomized controlled trial characteristics.
@@
We implement a “co-effect” model in which $T$ shifts both $Y_{1}$ and $Y_{2}$ with arm-specific probabilities and $W$ modifies only $Y_{1}$:
$\Pr(Y_{2}{=}1\mid T{=}1)=p^{(Z)}_{1},\quad\Pr(Y_{2}{=}1\mid T{=}0)=p^{(Z)}_{0}$,
@@
$p^{(Z)}_{1}=0.5+\tfrac{\mathrm{ATE}}{2},\quad p^{(Z)}_{0}=0.5-\tfrac{\mathrm{ATE}}{2},$
@@
$p^{(Y)}_{11}=0.5+\tfrac{\mathrm{ATE}}{2},\quad p^{(Y)}_{01}=0.5-\tfrac{\mathrm{ATE}}{2},\quad p^{(Y)}_{10}=0.2+\mathrm{ATE},\quad p^{(Y)}_{00}=0.2.$
@@
FM features. Each image $x$ is encoded with SigLIP [Zhai et al., 2023] into a patch-level representation; we use the final-layer token features (dim d=768d{=}768, $196$ patches/token positions). Unless noted otherwise, we do not use any task-specific fine-tuning.
@@ D.2 ISTAnt
We considered the randomized controlled trial introduced by Cadei et al. [2024]. Videos of ant triplets were collected under randomized treatment/control assignment. Throughout our unsupervised pipeline, domain annotations from biologists were used only a posteriori for interpretation/evaluation of discovered codes, never for training, as discussed in the main text.
FM features. Each frame $X$ is encoded with DINOv2 [Oquab et al., 2023] into a patch-level representation; we use the final-layer token features (dim d=384d{=}384, $256$ patches/token positions). Unless noted otherwise, we do not use any task-specific fine-tuning.
@@ Appendix E Additional Experiments
E.1 Evaluation on CelebA: what does our ground truth model?
We assess how well SAE codes behave as measurement channels on CelebA by aligning individual neurons with ground–truth attributes (see Section D.1). For each code $j$, we treat the event $Z_{j}>0$ as a binary predictor and compute its F1–score against the attribute label. The two most predictive neurons for the two affected factors are: (i) neuron 38 for Wearing_Hat with $\mathrm{F1}=0.841$, and (ii) neuron 6051 for Eyeglasses with $\mathrm{F1}=0.748$. Qualitative inspection of the top–activated images (Figure 6) confirms that these codes fire on the intended visual concept, supporting their use for exploratory causal inference.
@@ E.1 Evaluation on CelebA: what does our ground truth model?
Figure 6: Qualitative neurons’ interpretations. Each panel shows the 12 most–activated test images for the most predictive neuron of each affected outcome concept (activation = highest code value).
At the same time, the F1–score spectra over all neurons reveal a familiar pattern: a single, dominant “monosemantic” code per concept, accompanied by a long tail of weaker yet clearly non–zero correlations (Figure 7). This tail is stronger for Eyeglasses, where several neurons reach moderate F1, indicating broader leakage/entanglement. As discussed in the main text (see Section 3), such low–amplitude but widespread correlations are precisely what trigger the Paradox of Exploratory Causal Inference: with enough power, standard multi–testing will flag all of these leakage neurons as “significant.” Our NES counters this by retrieving the leading effect first and then recursively stratifying on previously discovered codes, so that subsequent tests target the residual causal signal rather than its leakage.
Figure 7: Monosemantic peaks with entanglement tails. For each attribute, we rank SAE codes by F1 against the CelebA label and visualize the top performers in order.
E.2 In–depth analysis: full semi–synthetic results
This subsection expands the quantitative picture in Figure 4 by showing the more complete grid of results across sample size and effect magnitude, for two evaluation regimes:

### v2

@@ Appendix B Neural Effect Test
2:  // A) Arm-wise residualize only the tested neuron $j$ (optional)
@@ Appendix C Minimal Python Implementation Snippet
We provide a minimal Python snippet for our algorithm relying on pandas library [mckinney2011pandas] for tabular operations and SciPy library for statistical testing [virtanen2020scipy] .
@@
By stratification with median binning. For simplicity, we ignore here arm-wise residualization (see full code for detailed implementation).
@@ D.1 CelebA semi-synthetic RCTs
Dataset. We use CelebA [liu2018large], a face attributes dataset with $200\mathrm{k}$ images and $40$ binary attributes per image  [^2] It can be downloaded from flwrlabs/celeba.. Furthermore, for implementation details, labels have been doubled (we pass from Beard to Has_Beard and Has_notBeard). We follow the authors’ official train/val/test split, and we employ the validation data for training SAEs and the test data to interpret them. Attributes are treated as ground-truth binary labels. From this source, we simulate several RCTs following the data-generating process (DGP) described below, varying the sample size ($n\ll 200k$) and treatment effect ($\tau$), reflecting realistic randomized controlled trial characteristics.
@@
We implement a “co-effect” model in which an intervention on $T$ shifts both $Y_{1}$ and $Y_{2}$ by the same ATE with arm-specific probabilities and $W$ modifies only $Y_{1}$, i.e.,:
$\Pr(Y_{2}{=}1\mid T{=}1)=p^{(Y_{2})}_{1},\quad\Pr(Y_{2}{=}1\mid T{=}0)=p^{(Y_{2})}_{0}$,
@@
$p^{(Y_{2})}_{1}=0.5+\tfrac{\mathrm{ATE}}{2},\quad p^{(Y_{2})}_{0}=0.5-\tfrac{\mathrm{ATE}}{2},$
@@
$p^{(Y_{1})}_{11}=0.5+\tfrac{\mathrm{ATE}}{2},\quad p^{(Y_{1})}_{01}=0.5-\tfrac{\mathrm{ATE}}{2},\quad p^{(Y_{1})}_{10}=0.2+\mathrm{ATE},\quad p^{(Y_{1})}_{00}=0.2.$
@@
FM features. Each image $x$ is encoded with SigLIP [zhai2023sigmoid] into a patch-level representation; we use the final-layer token features (dim d=768d{=}768, $196$ patches/token positions averaged). Unless noted otherwise, we do not use any task-specific fine-tuning.
@@
For each trained representation, including the additional experiments in Appendix E, we repeat the evaluation resampling the experiment 10 times and reporting the average $\pm$ standard deviation. Further ablations additionally retraining the encoder, i.e., SAE, are presented in Appendix E.
@@ D.2 ISTAnt
We considered the randomized controlled trial introduced by cadei2024smoke. Videos of ant triplets were collected under randomized treatment/control assignment. Throughout our unsupervised pipeline, domain annotations from biologists were used only a posteriori for interpretation/evaluation of discovered codes, never for training, as discussed in the main text.
FM features. Each frame $X$ is encoded with DINOv2 [oquab2023dinov2] into a patch-level representation; we use the final-layer token features (dim d=384d{=}384, $256$ patches/token positions averaged). Unless noted otherwise, we do not use any task-specific fine-tuning.
@@ Appendix E Additional Experiments
In this Section, we provide extensive empirical validation of our method, in addition to the selected main semi-synthetic experiments reported in Section 6. Particularly,
  • we explicitly test for the required encoder representation assumptions, theoretically justifying the successful identification results;
  • we extensively validate the method consistency by (i) repeating the experiments with different encoders and trainings, (ii) comparing different method hyper-parameters, and (iii) additionally testing on more varied data-generating processes;
  • we additionally include two novel heuristic baselines, inspired by Causal Feature Learning principles [chalupka2017causal].
Overall, the experiments consistently validate our identification result and further provide practical guidelines on how to set the different hyperparameters.
E.1 Assumptions Validation
We assess how well SAE codes behave as measurement channels on CelebA by testing the alignment between individual neurons and the ground–truth attributes described in the previous section. We consider the main SAE trained for the main experiments reported in Section 6 with training details described in Appendix D, and we then compare it with the same encoder replacing SigLIP (foundational model) with DINOv2, and again SAE top-$k$ non linearity with JumpRELU. According to Theorem 4.1, we test if (i) such attributes information is still encoded in the SAE representation, i.e., Assumption A.1 and (ii) per each attribute there exists a dedicated neuron strongly aligned and all the other neurons are just marginally aligned with it, i.e., Assumption A.2. We first test the foundational model sufficiency and then for the SAE. Particularly, for each code $j$ in the SAE representation, we treat the event $Z_{j}>0$ as a binary predictor and compute its F1–score against each attribute label (assumed given). We then interpret the most predictive neurons by visualizing their corresponding most activated samples; and visualize such attributes’ entanglement in the representation by filtering all their most predictive codes and comparing their probing performances. We remark that in practice, without the affected outcome supervision, these assumptions are not testable, and we can do it here just because controlling the data-generating process and having access to attribute annotations. More indirect and unsupervised assessments for a SAE representation, testing a priori validity for NES, should be explored in future works, or just ignored, at the price of having no guarantees on the interpretability of NES results.
SigLIP with SAE top-$k$ non linearity
(main encoder). As expected, SigLIP embeddings are sufficient for the task, enabling us to accurately classify both the affected outcomes directly via (vanilla) logistic regression, respectively with Precision=0.9639=0.9639 and Recall=0.9947=0.9947 for Wearing_Hat, and with Precision=0.9712=0.9712 and Recall=0.9747=0.9747 for Eyeglasses. Furthermore, in the SAE representation, trained with top-$k$ non-linearity ($k=20$), the two most predictive neurons for the two affected factors are: (i) neuron 38 for Wearing_Hat with $\mathrm{F1-score}=0.841$, and (ii) neuron 6051 for Eyeglasses with $\mathrm{F1-score}=0.748$. We can further appreciate qualitatively the high separations by visualizing the most activated images per neuron in Figure 7. Also, the SAE representation is then sufficient, and for each affected outcome, there is a principal neuron strongly aligned. In Figure 8 we visualize the F1-score spectra of the most predictive neurons for each attribute. The long tail, dominated by a principal component, represents exactly what is already required by Assumptions A.1-A.2 to enforce NES validity. Such low–amplitude but widespread correlations are precisely what triggers instead the Paradox of Exploratory Causal Inference with vanilla multiple-testing, i.e., with enough power all such entangled neurons are returned as significantly affected due to their (marginal) true affected outcome information. Our NES counters this issue by retrieving the leading effect first and then recursively stratifying on previously discovered codes, so that subsequent tests target the residual causal signal rather than previous causes’ information leakages.
@@ E.1 Assumptions Validation
Figure 7: Qualitative neurons’ interpretations (SigLIP + SAE top-$k$). Each panel shows the 12 most–activated test images for the most predictive neuron of each affected outcome concept (activation = highest code value). Both concepts are perfectly separated.
Figure 8: Assessing principal alignment with broad entanglement (SigLIP + SAE top-$k$). For each attribute, we rank SAE codes by F1-score against the affected attributes and visualize the most predictive in order. Both concepts are principally explained by a single neuron, and still weakly, while broadly entangled with many others.
DINOv2 with SAE top-$k$ non-linearity
As expected, also DINOv2 embeddings are sufficient for the task, enabling to accurately classify both the affected outcomes directly via (vanilla) logistic regression, respectively with Precision=0.9534=0.9534 and Recall=0.9787=0.9787 for Wearing_Hat, and with Precision=0.9636=0.9636 and Recall=0.9567=0.9567 for Eyeglasses. Furthermore, in the SAE representation, trained with top-$k$ non-linearity ($k=20$), the two most predictive neurons for the two affected factors are: (i) neuron 1937 for Wearing_Hat with $\mathrm{F1-score}=0.431$, and (ii) neuron 253 for Eyeglasses with $\mathrm{F1-score}=0.594$. We can further appreciate qualitatively the separations by visualizing the most activated images per neuron in Figure 9. Also, the SAE representation is then sufficient, while the principal alignment is high for Eyeglasses but less evident for Wearing_Hat, where the concepts could be possibly captured by all the top three most predictive neurons. Indeed, also in the qualitative interpretation, among the most activated samples, some concepts, e.g., wearing sunglasses and a hat, are over-represented. In Figure 10, we visualize the F1-score spectra of the most predictive neurons for each attribute. The long tail, dominated by a principal component, represents exactly what is already required by Assumptions A.1-A.2 to enforce NES validity, even if the principal alignment is weakly satisfied by Wearing_Hat concept (potentially violating Equation 26). Such low–amplitude but widespread correlations are precisely what trigger the Paradox of Exploratory Causal Inference with vanilla multiple-testing, i.e., with enough power, all such entangled neurons are returned as significantly affected due to their (marginal) true affected outcome information. Our NES counters this issue by retrieving the leading effect first and then recursively stratifying on previously discovered codes, so that subsequent tests target the residual causal signal rather than previous causes’ information leakages.
Figure 9: Qualitative neurons’ interpretations (DINOv2 + SAE top-$k$). Each panel shows the 12 most–activated test images for the most predictive neuron of each affected outcome concept (activation = highest code value). Both concepts are perfectly separated, despite Wearing_Hat seeming to over-represent specific subgroups.
Figure 10: Assessing principal alignment with broad entanglement (DINOv2 + SAE top-$k$). For each attribute, we rank SAE codes by F1-score against the affected attributes and visualize the most predictive in order. Eyeglasses is principally explained by a single neuron, Wearing_Hat by three main neurons, and both concepts are still weakly but broadly entangled with many others.
SigLIP with SAE JumpReLU non-linearity
We already discussed SigLIP embedding sufficiency in the main encoder hypotheses assessment. However, training on top a SAE with jump-ReLU non-linearity breaks the desiderata hypotheses satisfied with top-$k$ non-linearity. Particularly, the two most predictive neurons for the two affected factors are: (i) neuron 3485 for Wearing_Hat with $\mathrm{F1-score}=0.273$, and (ii) neuron 2865 for Eyeglasses with $\mathrm{F1-score}=0.425$. Both performances are limited, particularly for Wearing_Hat classification. We can further appreciate qualitatively the weak separability by visualizing the most activated images per neuron in Figure 11, e.g., Neuron 3485 is not capturing precisely the concept Wearing_Hat but probably something broader like “having something on top of the head”, e.g., some text from the background. In Figure 10, we visualize the F1-score spectra of the most predictive neurons for each attribute. For Eyeglasses, there exists a principal neuron, despite the other entangled neurons still being sufficiently predictive. Instead, for Wearing_Hat classification, there is no clear principal component, violating NES assumption, and also ill-defining any evaluation attempts (requiring setting the principal neurons identifying the effects). On such a representation, NES has no guarantees, and the long and heavy tails also trigger once again the Paradox of Exploratory Causal Inference with vanilla multiple-testing.
Figure 11: Qualitative neurons’ interpretations (SigLIP + SAE jump-ReLU). Each panel shows the 12 most–activated test images for the most predictive neuron of each affected outcome concept (activation = highest code value). Eyeglasses is separated while it is not the case for Wearing_Hat concept.
Figure 12: Assessing principal alignment with broad entanglement (SigLIP + SAE jump-ReLU). For each attribute, we rank SAE codes by F1-score against the affected attributes and visualize the most predictive in order. Eyeglasses concept is principally explained by a single neuron, and still strongly and broadly entangled with many others. Wearing_Hat concept has no principal components and the tails are very heavy, i.e., the information is shared among different neurons.
E.2 Method Consistency
In this Section, we extensively validate the method consistency beyond the experiments selected for the main results discussion in Section 6. Particularly, (i) we repeat the experiments with different encoders and trainings, (ii) we compare different method hyper-parameters, and (iii) we additionally test on more varied data-generating processes.
E.2.1 Varying Encoder
Considering the same data-generating processes and evaluations described in Appendix D, we repeat the experiments relying on different measurement channels, i.e., input encoders and corresponding embedding.
Foundational model
First, we replicate the main experiment replacing SigLIP with DINOv2 [oquab2023dinov2] as foundational model. From the experiment in the previous paragraph, we know that the main hypotheses for NES to work hold, still triggering the Paradox of Causal Inference if relying otherwise on vanilla multiple hypothesis testing. The results are reported in Figure 13. As expected, DINOv2 perfectly replicates the results discussed in Section 6.
Figure 13: Ablation Foundationa Model (DINOv2). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), relying on DINOv2 embeddings. NES consistently replicates the main results using SigLIP embeddings, still overcoming the paradox of Exploratory Causal Inference.
Sparse autoencoder dimension
Then, we consider different sparse autoencoders varying the latent dimension $m$, and retrying with the same setting described in Appendix D. Particularly, we varied $m$ by a factor $\{4,6,8,10,14,16\}$ of the foundational model dimension $d=768$ (in the main experiment, we considered $m=12\cdot d$). In all the settings and retraining, we replicate the results discussed in Section 6, i.e., NES consistently overcomes the Paradox of Exploratory Causal Inference, while it consistently invalidates all the other baselines, making both our method and the baselines invariant with respect to the latent representation dimension, in the order of a few thousand neurons. See results summaries in Figures 14-19.
Figure 14: Ablation SAE dimension (m=3 072m=3\ 072). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE dimension m=3 072m=3\ 072, 44x DINOv2 dimension $n=768$. NES consistently replicates the main results, varying the representation dimension $m$, still overcoming the paradox of Exploratory Causal Inference.
Figure 15: Ablation SAE dimension (m=4 608m=4\ 608). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE dimension m=4 608m=4\ 608, 66x DINOv2 dimension $n=768$. NES consistently replicates the main results, varying the representation dimension $m$, still overcoming the paradox of Exploratory Causal Inference.
Figure 16: Ablation SAE dimension (m=6 144m=6\ 144). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE dimension m=6 144m=6\ 144, 88x DINOv2 dimension $n=768$. NES consistently replicates the main results, varying the representation dimension $m$, still overcoming the paradox of Exploratory Causal Inference.
Figure 17: Ablation SAE dimension (m=7 680m=7\ 680). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE dimension m=7 680m=7\ 680, 1010x DINOv2 dimension $n=768$. NES consistently replicates the main results, varying the representation dimension $m$, still overcoming the paradox of Exploratory Causal Inference.
Figure 18: Ablation SAE dimension (m=10 752m=10\ 752). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE dimension m=10 752m=10\ 752, 1414x DINOv2 dimension $n=768$. NES consistently replicates the main results, varying the representation dimension $m$, still overcoming the paradox of Exploratory Causal Inference.
Figure 19: Ablation SAE dimension (m=12 288m=12\ 288). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE dimension m=12 288m=12\ 288, 1616x DINOv2 dimension $n=768$. NES consistently replicates the main results, varying the representation dimension $m$, still overcoming the paradox of Exploratory Causal Inference.
Sparse autoencoder non-linearity
Additionally, we consider different sparse auto-encoder non-linearities and repeat all the experiments as before. In Figures 20-23 we report all the results using top-$k$ activation with $k\in\{5,10,50,100\}$ (by default $k=20$ in the main experiments). Again, all the results discussed in Section 6 are replicated, i.e., NES consistently overcomes the Paradox of Exploratory Causal Inference, suggesting the top-$k$ selection is not influential for our method, i.e., the required assumptions are consistently satisfied, despite the obtained representation may significantly change. On the other hand, the other baselines consistently fail due to the Paradox of Exploratory Causal Inference, and the higher the enforced entanglement, i.e., $k$, the higher the broad entanglement with the affected outcomes, and accordingly smaller the precision. Finally, any analysis using jump-ReLU for non-linearity is invalid due to the principal alignment discussed above, required both for NES validity and more generally for valid evaluation. For completeness, we report anyway in Figure 24 the replicated experiments results using jump-ReLU non-linearity. At least one, if not both, ground truths are ill-defined, and the consistent null precision, even for NES, should be taken with a pinch of salt.
Figure 20: Ablation SAE non-linearity (top-55). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE non-linearity by top-$k$ with $k=5$. NES consistently replicates the main results using top-$k$ with $k=20$, still overcoming the paradox of Exploratory Causal Inference.
Figure 21: Ablation SAE non-linearity (top-$10$). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE non-linearity by top-$k$ with $k=10$. NES consistently replicates the main results using top-$k$ with $k=20$, still overcoming the paradox of Exploratory Causal Inference.
Figure 22: Ablation SAE non-linearity (top-$50$). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE non-linearity by top-$k$ with $k=50$. NES consistently replicates the main results using top-$k$ with $k=20$, still overcoming the paradox of Exploratory Causal Inference.
Figure 23: Ablation SAE non-linearity (top-$100$). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE non-linearity by top-$k$ with $k=100$. NES consistently replicates the main results using top-$k$ with $k=20$, still overcoming the paradox of Exploratory Causal Inference.
Figure 24: Ablation SAE non-linearity (jump-ReLU). Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), with SAE non-linearity by jump-ReLU. The principal alignment is violated, invalidating both NES and any evaluation (ill-defined ground truth).
Seed
Finally, we replicate two more times exactly the same main experiments presented in Section 6), retraining the SAE with a different random seed. Once again, we replicate the results discussed in Section 6, i.e., NES consistently overcomes the Paradox of Exploratory Causal Inference, while it invalidates all the other baselines. See results in Figures 25-26.
Figure 25: Replica-1 main results. Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), retraining the SAE with a different random seed (replica 1). NES consistently replicates the main results, still overcoming the paradox of Exploratory Causal Inference.
Figure 26: Replica-2 main results. Precision, Recall, and IoU in effect identification varying sample size NN (top) and effect size $\tau$ (bottom), retraining the SAE with a different random seed (replica 2). NES consistently replicates the main results, still overcoming the paradox of Exploratory Causal Inference.
E.2.2 Varying Method hyper-parameters
We now repeat the main experiments and compare different variants of our methods, varying the multiple-hypothesis testing, using residualization and varying the ATE estimator.
Hypothesis Testing
[…]
Figure 27: Ablation multi-hypotheses testing (within NES). Bonferroni and FDR: best recoveries, full recall, keeping Precision $\gtrsim 0$ with sufficient experiment power; $t$-test: comparable overall, with similar performances with higher power, potentially more exploratory for very small power, and less precise in mid-power regimes.
Residualization
We then evaluate the benefits in combining the recursive stratification procedure characterizing NES, with or without arm-wise residualization within NeuralEffectTest (lines 2-10) for each multi-hypotheses testing method. Same experiment set-up as described in Appendix D. Results are reported in Figure 28. In lower power experiments, i.e., smaller $n$ or $\tau$, arm-wise residualization seems to consistently improve the precision without compromising the recall, for each multi-hypotheses testing variant. In higher power experiments, i.e., greater $n$ or $\tau$, such difference gets negligible as discussed in Theorem 4.1 discussion in Appendix A.
Figure 28: Ablation arm-wise residualization (within NET). Additional arm-wise residualization significantly improve the method performances in lower power regime, i.e., smaller sample or reduced effect.
Average Treatment Effect estimator
Throughout the paper, our per-neuron hypothesis test uses the associational difference (AD), i.e., a two-sample $t$-test on the treated–control difference in means. In randomized trials, AD is unbiased for the ATE, but it is not semi-parametrically efficient. A standard variance–reduction alternative is Augmented Inverse Propensity Weighting (AIPW; robins1994estimation), which orthogonalizes the estimator against misspecification of either the propensity score or the outcome regression. For each code $j$, let $Z_{ij}$ be its activation for unit $i$, $T_{i}\!\in\!\{0,1\}$ the treatment, and $W_{i}$ observed exogenous causes. We compute the AIPW pseudo-outcome
where $\pi(W)\!=\!\Pr(T\!=\!1\mid W)$, known and constant $\pi=0.5$ in balanced RCTs, and $\hat{\mu}_{tj}(W)\approx\mathbb{E}[Z_{j}\mid T\!=\!t,W]$ is a nuisance regression. The AIPW estimate of the code-level ATE is $\hat{\tau}^{\text{AIPW}}_{j}=\frac{1}{n}\sum_{i}\tilde{Z}_{ij}$; we test $H_{0}:\tau_{j}=0$ via a one-sample $t$-test on $\{\tilde{Z}_{ij}\}_{i}$ with robust variance. Figure 29 reports the standalone multiple testing methods performances comparing AD vs. AIPW as neural treatment effect estimator, fixing the effect magnitude $\tau=0.6$ and across sample sizes $n\in\{30,500,5\ 000\}$. In our setting, with balanced and randomized treatment assignment ($\pi=0.5$) and a single binary effect-modifier $W$, AIPW yields only marginal efficiency gains: Precision/Recall/IoU curves are essentially overlapping, with small stability improvements for AIPW at the smallest $n$. Crucially, orthogonalization affects variance but does not resolve entanglement: the significance–collapse phenomenon for standard multi-testing (Section 3) persists under AIPW, and NES retains its advantage because its benefit comes from recursive stratification (disentangling residual effects), not from how the first-step mean contrast is estimated. Takeaways:
  • in pure RCTs with weak, low-dimensional $W$, AD is competitive and simpler,
  • AIPW can be preferred when richer exogenous information is available (higher-dimensional $W$, imbalance, or mild protocol deviations), where its variance reduction can translate into earlier detection of the leading effect,
  • regardless of AD or AIPW, NES’s stratified recursion is the key to avoiding over-discovery under entanglement.
Figure 29: Ablation Treatment Effect Estimator. Precision, Recall, and IoU when replacing the per-neuron causal estimator AD (associational difference) with AIPW for vanilla multiple testing criteria (baseline).
E.2.3 Varying Structural Causal Model
Additional main results
This paragraph expands the quantitative results illustrated in Figure 5 by reporting all the considered sample sizes and effect magnitudes in the data-generating processes, for two evaluation regimes: